# Gamma-Ray Observations of Galaxy Clusters Strongly Constrain Dark Matter Annihilation in Prompt Cusps

### v1

@@ I Introduction
In standard cosmological models, “prompt cusps” of high dark matter density may efficiently form during certain phases in the early Universe Diemand et al. (2005); Ishiyama et al. (2010); Anderhalden and Diemand (2013); Ishiyama (2014); Polisensky and Ricotti (2015); Ogiya et al. (2016); Angulo et al. (2017); Delos et al. (2018a, b); Ogiya and Hahn (2018); Ishiyama and Ando (2020); Colombi (2021) — a finding recently supported by novel analytical calculations White (2022); Del Popolo and Fakhry (2023). Intriguingly, studies indicate that these cusps, which follow a $\rho_{\text{cusp}}\propto r^{-1.5}$ radial density profile, form almost instantaneously Delos et al. (2019a); Delos and White (2022); Ondaro-Mallea et al. (2023) and survive both halo growth Delos and White (2022) and clustering Delos and White (2023).
@@
The similarity between annihilation in prompt cusps and dark matter decay motivates $\gamma$-ray observations of galaxy clusters. As the most massive structures in the Universe, galaxy clusters stand among the best targets for the indirect detection of dark matter decay Ackermann et al. (2010); Huang et al. (2012); Ando and Nagai (2012); Ackermann et al. (2015a); Tan et al. (2020); Thorpe-Morgan et al. (2021); Di Mauro et al. (2023); Song et al. (2024). Notably, galaxy cluster constraints on decaying dark matter can set lower-limits on the dark matter lifetime that exceed 1027 s Song et al. (2024).
@@
Our analysis leverages the enhanced annihilation rate from prompt cusps to improve the sensitivity of $\gamma$-ray searches for dark matter. Most studies focus on $\gamma$-ray fluxes that scale with the square of dark matter density, favoring targets with high central densities, such as the Galactic Center Hooper and Goodenough (2011); Hooper and Linden (2011); Abazajian and Kaplinghat (2012); Daylan et al. (2016); Calore et al. (2015b, a); Ackermann et al. (2017); Leane and Slatyer (2019); Di Mauro (2021) or dwarf spheroidal galaxies (dSphs)  Abdo et al. (2010); Ackermann et al. (2014); Geringer-Sameth et al. (2015a, b); Ackermann et al. (2015b); Li et al. (2016); Albert et al. (2017); Chiappo et al. (2019); Hoof et al. (2020); Li et al. (2018); Calore et al. (2018); Alvarez et al. (2020); Linden (2020); McDaniel et al. (2024); Li et al. (2021); Crnogorčević and Linden (2024). However, the annihilation rate from prompt cusps linearly depends on the total dark matter mass, elevating high mass targets (e.g., galaxy clusters or the isotropic $\gamma$-ray background (IGRB) Delos et al. (2024)), to the most promising candidates.
@@ II.1 Cosmological distribution of prompt cusps
These prompt cusps grow halos around them over time, and the halos also contribute to dark matter annihilation. However, prior halo and subhalo modeling suggests that their contribution to the annihilation rate is only at about the 10 percent level Delos and White (2023), so we neglect them.
@@
$\rho_{\mathrm{eff},0}\equiv\frac{\int_{\mathrm{cusps}}\rho^{2}\mathrm{d}V}{M}=n_{\mathrm{peaks}}\langle j\rangle/\bar{\rho}_{0}.$ (6)
For a dark matter particle of mass $m_{\mathrm{DM}}$ and cross section $\langle\sigma v\rangle$ (which we assume to be velocity-independent at lowest order), the cusps would contribute an average annihilation rate
$\left\langle\frac{\Gamma}{M}\right\rangle_{\!0}$ $=\frac{\langle\sigma v\rangle}{2m_{\mathrm{DM}}^{2}}\rho_{\mathrm{eff},0}$ (7)
@@ II.2 Annihilation in galaxy clusters
$\frac{\mathrm{d}^{2}\Phi}{\mathrm{d}\Omega\mathrm{d}E}$ $=\frac{1}{4\pi}\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}E}\int_{0}^{\infty}\mathrm{d}\ell\,\rho(r)\left\langle\frac{\Gamma}{M}\right\rangle,$ (8)
given in terms of the annihilation rate per mass of dark matter, $\langle\Gamma/M\rangle$. Here, $\ell$ is the position along the observer’s line of sight, and the cluster-centric radius $r$ accordingly depends on $\ell$. We expand the annihilation rate as
$\left\langle\frac{\Gamma}{M}\right\rangle$ $=\frac{\langle\sigma v\rangle}{2m_{\mathrm{DM}}^{2}}\left[\rho(r)+f_{\mathrm{surv}}f_{\mathrm{tidal}}(r)\rho_{\mathrm{eff},0}\right].$ (9)
The first term represents the smooth halo, which contains almost all of the dark matter, as only about 1% resides in prompt cusps Delos and White (2023). The second term accounts for the cusps and is constructed by scaling Eq. (7) by the cusp survival factor $f_{\mathrm{surv}}$ and by a position-dependent factor $f_{\mathrm{tidal}}(r)$.
Name RA, Dec (deg.) $J$ (GeV2 cm-5) distance
Virgo 187.70, 12.34 $8.88\times 10^{20}$ $17.053\pm 0.360$
NGC 4636 190.71, 2.69 $6.85\times 10^{19}$ $16.303\pm 0.578$
Hydra 139.53, $-12.09$ $4.99\times 10^{19}$ $245.41\pm 17.25$
Coma 194.95, 27.98 $4.95\times 10^{19}$ $109.350\pm 8.261$
M49 187.44, 8.00 $4.88\times 10^{19}$ $16.060\pm 0.348$
Centaurus 192.20, $-41.31$ $4.61\times 10^{19}$ $40.100\pm 2.177$
Fornax 54.62, $-35.45$ $4.12\times 10^{19}$ $17.711\pm 0.395$
Table 1: Clusters considered in our analysis. The annihilation $J$ factor depends mildly on the dark matter mass and decoupling temperature; here we fix $m_{\mathrm{DM}}=100$ GeV and $T_{\mathrm{d}}=30$ MeV.
@@
$\frac{\mathrm{d}J}{\mathrm{d}\Omega}\equiv\int_{0}^{\infty}\mathrm{d}\ell\,\rho(r)[\rho(r)+f_{\mathrm{surv}}f_{\mathrm{tidal}}(r)\rho_{\mathrm{eff},0}].$ (12)
@@
Table 1 gives the annihilation $J$ factor of each cluster for an example model with $m_{\mathrm{DM}}=100$ GeV and $T_{\mathrm{d}}=30$ MeV. Figure 2 shows the annihilation morphology over the sky. Due to its proximity, the Virgo cluster has the highest $J$ by far, but it also produces the most spatially extended signal.
Generally, higher dark matter masses and earlier decoupling (higher $T_{\mathrm{d}}$) both lead to larger values of $J$, because they cause the prompt cusps to be slightly more internally dense. However, $J$ depends only logarithmically on $m_{\mathrm{DM}}$ and $T_{\mathrm{d}}$ Delos and White (2023). We follow previous studies Ackermann et al. (2015c); Delos et al. (2024) in fixing $T_{\mathrm{d}}=30$ MeV for simplicity. This is a mildly conservative choice; lower $T_{\mathrm{d}}$ tend to be in tension with terrestrial experiments Cornell et al. (2013), since they require a strong kinetic coupling between the dark matter and the Standard Model. We however analyze a range of dark matter masses $m_{\mathrm{DM}}$, and Fig. 3 shows how each cluster’s $J$ factor varies with $m_{\mathrm{DM}}$ (with $T_{\mathrm{d}}=30$ MeV fixed). The spatial distribution of the annihilation signal also depends in principle on $m_{\mathrm{DM}}$, since denser cusps are less susceptible to tidal stripping. However, this variation is small enough that Fig. 2 does not discernibly change over the full range of $5~{}\text{GeV}<m_{\mathrm{DM}}<10^{5}~{}\text{GeV}$.
@@
Figure 4: Comparing different contributions to the annihilation signal of the Virgo cluster. As a function of the angle $\theta$ from the cluster center, we show the contribution $\mathrm{d}J/\mathrm{d}\ln\theta=2\pi\theta^{2}\mathrm{d}J/\mathrm{d}\Omega$ of each logarithmic angular interval to the integrated $J$ factor. The light teal solid curve shows the net predicted $J$ factor, which we use for our analysis. The navy dotted curve shows the $J$ factor without accounting for tidal stripping of prompt cusps (5–10% higher). The magenta dashed curve shows the $J$ factor from the smooth halo alone, neglecting contributions from cusps altogether.

### v2

@@ I Introduction
In standard cosmological models, “prompt cusps” of high dark matter density may efficiently form during certain phases in the early Universe Diemand et al. (2005); Ishiyama et al. (2010); Anderhalden and Diemand (2013); Ishiyama (2014); Polisensky and Ricotti (2015); Ogiya et al. (2016); Angulo et al. (2017); Delos et al. (2018a, b); Ogiya and Hahn (2018); Ishiyama and Ando (2020); Colombi (2021) — a finding recently supported by novel analytical calculations White (2022); Del Popolo and Fakhry (2023). Intriguingly, studies indicate that these cusps, which follow a $\rho_{\text{cusp}}\propto r^{-1.5}$ radial density profile, form almost instantaneously Delos et al. (2019a); Delos and White (2022); Ondaro-Mallea et al. (2023) and survive both halo growth Delos and White (2022) and clustering Delos and White (2023); Delos (2025).
@@
The similarity between annihilation in prompt cusps and dark matter decay motivates $\gamma$-ray observations of galaxy clusters. As the most massive structures in the Universe, galaxy clusters stand among the best targets for the indirect detection of dark matter decay Ackermann et al. (2010); Huang et al. (2012); Ando and Nagai (2012); Tan et al. (2020); Thorpe-Morgan et al. (2021); Di Mauro et al. (2023); Song et al. (2024a). Notably, galaxy cluster constraints on decaying dark matter can set lower-limits on the dark matter lifetime that exceed 1027 s Song et al. (2024a).
@@
Our analysis leverages the enhanced annihilation rate from prompt cusps to improve the sensitivity of $\gamma$-ray searches for dark matter. Most studies focus on $\gamma$-ray fluxes that scale with the square of dark matter density, favoring targets with high central densities, such as the Galactic Center Hooper and Goodenough (2011); Hooper and Linden (2011); Abazajian and Kaplinghat (2012); Zhou et al. (2015); Daylan et al. (2016); Calore et al. (2015b, a); Ackermann et al. (2017); Leane and Slatyer (2019); Abazajian et al. (2020); Di Mauro (2021); Song et al. (2024b) or dwarf spheroidal galaxies (dSphs)  Abdo et al. (2010); Ackermann et al. (2014); Geringer-Sameth et al. (2015a, b); Ackermann et al. (2015a); Li et al. (2016); Albert et al. (2017); Chiappo et al. (2019); Hoof et al. (2020); Li et al. (2018); Calore et al. (2018); Alvarez et al. (2020); Linden (2020); McDaniel et al. (2024); Li et al. (2021); Crnogorčević and Linden (2024). However, the annihilation rate from prompt cusps linearly depends on the total dark matter mass, elevating high mass targets (e.g., galaxy clusters or the isotropic $\gamma$-ray background (IGRB) Delos et al. (2024)), to the most promising candidates.
@@ II.1 Cosmological distribution of prompt cusps
These prompt cusps grow halos around them over time, and the halos also contribute to dark matter annihilation. However, prior halo and subhalo modeling suggests that their contribution to the annihilation rate is only at about the 10 percent level Delos and White (2023), so we neglect them. We will revisit this comparison in Section II.2.
@@
$\langle\rho_{\mathrm{eff}}\rangle_{0}\equiv\frac{\int_{\mathrm{cusps}}\rho^{2}\mathrm{d}V}{M}=n_{\mathrm{peaks}}\langle j\rangle/\bar{\rho}_{0}.$ (6)
For a dark matter particle of mass $m_{\mathrm{DM}}$ and velocity-weighted annihilation cross-section $\langle\sigma v\rangle$ (which we assume to be velocity-independent at lowest order), the cusps would contribute an average annihilation rate
$\left\langle\frac{\Gamma}{M}\right\rangle_{\!0}$ $=\frac{\langle\sigma v\rangle}{2m_{\mathrm{DM}}^{2}}\langle\rho_{\mathrm{eff}}\rangle_{0}$ (7)
@@ II.2 Annihilation in galaxy clusters
$\frac{\mathrm{d}^{2}\Phi}{\mathrm{d}\Omega\mathrm{d}E}$ $=\frac{1}{4\pi}\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}E}\int_{\text{l.o.s.}}\mathrm{d}\ell\,\rho(r)\left\langle\frac{\Gamma}{M}\right\rangle,$ (8)
given in terms of the annihilation rate per mass of dark matter, $\langle\Gamma/M\rangle$. Here, $\ell$ is the position along the observer’s line of sight, the cluster-centric radius $r$ accordingly depends on $\ell$, and the line-of-sight integration is performed out to the virial radius of each cluster. We expand the annihilation rate as
$\left\langle\frac{\Gamma}{M}\right\rangle$ $=\frac{\langle\sigma v\rangle}{2m_{\mathrm{DM}}^{2}}\left[\rho(r)+f_{\mathrm{surv}}f_{\mathrm{tidal}}(r)\langle\rho_{\mathrm{eff}}\rangle_{0}\right].$ (9)
The first term represents the smooth halo Salucci (2019), which contains almost all of the dark matter, as only about 1% resides in prompt cusps Delos and White (2023). The second term accounts for the cusps and is constructed by scaling Eq. (7) by the cusp survival factor $f_{\mathrm{surv}}$ and by a position-dependent factor $f_{\mathrm{tidal}}(r)$.
Name RA, Dec (deg.) distance (Mpc) $\rho_{s}$ (MMpc-3) $r_{\mathrm{s}}$ (Mpc) $R_{200}$ (Mpc) $\theta_{200}$ $M_{200}$ (M) $J$ (GeV2 cm-5)
Virgo 187.70, 12.34 $15.46$ $1.22\times 10^{15}$ $0.335$ $1.700$ $6.3^{\circ}$ $5.6\times 10^{14}$ $8.88\times 10^{20}$
NGC 4636 190.71, 2.69 $17.18$ $1.64\times 10^{15}$ $0.135$ $0.777$ $2.6^{\circ}$ $5.3\times 10^{13}$ $6.85\times 10^{19}$
Hydra 139.53, 12.09-12.09 $47.51$ $1.3\times 10^{15}$ $0.264$ $1.376$ $1.7^{\circ}$ $3.0\times 10^{14}$ $4.99\times 10^{19}$
Coma 194.95, 27.98 $100.24$ $1.17\times 10^{15}$ $0.454$ $2.260$ $1.3^{\circ}$ $1.3\times 10^{15}$ $4.95\times 10^{19}$
M49 187.44, 8.00 $18.91$ $1.68\times 10^{15}$ $0.127$ $0.741$ $2.2^{\circ}$ $4.6\times 10^{13}$ $4.88\times 10^{19}$
Centaurus 192.20, 41.31-41.31 $43.16$ $1.33\times 10^{15}$ $0.239$ $1.258$ $1.7^{\circ}$ $2.3\times 10^{14}$ $4.61\times 10^{19}$
Fornax 54.62, 35.45-35.45 $21.50$ $1.65\times 10^{15}$ $0.132$ $0.763$ $2.0^{\circ}$ $5.1\times 10^{13}$ $4.12\times 10^{19}$
Table 1: Clusters considered in our analysis and the properties that we adopt. For each cluster, we adopt a Navarro-Frenk-White (NFW) density profile, $\rho(r)=\rho_{\mathrm{s}}(r/r_{\mathrm{s}})^{-1}(1+r/r_{\mathrm{s}})^{-2}$ Navarro et al. (1996, 1997), extending out to the virial radius $R_{200}$, which encloses mass $M_{200}$ and corresponds to an angle $\theta_{200}$ on the sky. Distances and structural parameters are drawn from Ref. Di Mauro et al. (2023). We also show the integrated annihilation $J$ factor calculated in accordance with Section II. $J$ depends mildly on the dark matter mass and decoupling temperature; here we fix $m_{\mathrm{DM}}=100$ GeV and $T_{\mathrm{d}}=30$ MeV.
@@
$\frac{\mathrm{d}J}{\mathrm{d}\Omega}\equiv\int_{\text{l.o.s.}}\mathrm{d}\ell\,\rho(r)[\rho(r)+f_{\mathrm{surv}}f_{\mathrm{tidal}}(r)\langle\rho_{\mathrm{eff}}\rangle_{0}].$ (12)
@@
Table 1 gives the annihilation $J$ factor of each cluster (integrated out to the virial radius) for an example model with $m_{\mathrm{DM}}=100$ GeV and $T_{\mathrm{d}}=30$ MeV. Figure 2 shows the annihilation morphology over the sky. Due to its proximity, the Virgo cluster has the highest $J$ by far, but it also produces the most spatially extended signal.
Generally, higher dark matter masses and earlier decoupling (higher $T_{\mathrm{d}}$) both lead to larger values of $J$, because they cause the prompt cusps to be slightly more internally dense. However, $J$ depends only logarithmically on $m_{\mathrm{DM}}$ and $T_{\mathrm{d}}$ Delos and White (2023). We follow previous studies Ackermann et al. (2015b); Delos et al. (2024) in fixing $T_{\mathrm{d}}=30$ MeV for simplicity. This is a mildly conservative choice; lower $T_{\mathrm{d}}$ tend to be in tension with terrestrial experiments Cornell et al. (2013), since they require a strong kinetic coupling between the dark matter and the Standard Model. We however analyze a range of dark matter masses $m_{\mathrm{DM}}$, and Fig. 3 shows how each cluster’s $J$ factor varies with $m_{\mathrm{DM}}$ (with $T_{\mathrm{d}}=30$ MeV fixed). The spatial distribution of the annihilation signal also depends in principle on $m_{\mathrm{DM}}$, since denser cusps are less susceptible to tidal stripping. However, this variation is small enough that Fig. 2 does not discernibly change over the full range of $5~\text{GeV}<m_{\mathrm{DM}}<10^{5}~\text{GeV}$.
@@
Figure 4: Comparing different contributions to the annihilation signal of the Virgo cluster. As a function of the angle $\theta$ from the cluster center, we show the contribution $\mathrm{d}J/\mathrm{d}\ln\theta=2\pi\theta^{2}\mathrm{d}J/\mathrm{d}\Omega$ of each logarithmic angular interval to the integrated $J$ factor. The light teal solid curve shows the net predicted $J$ factor, which we use for our analysis. The navy dotted curve shows the $J$ factor without accounting for tidal stripping of prompt cusps (5–10% higher). The magenta dashed curve shows the $J$ factor from the smooth halo alone, neglecting contributions from cusps altogether. For comparison, the two gray lines show the contribution from subhalos (without prompt cusps) estimated according to two commonly used extrapolations (see the text).
Previous efforts to constrain dark matter annihilation using galaxy clusters and large-scale structure (e.g. Ackermann et al. (2015c); Chan and Leung (2017); Bartlett et al. (2022); Di Mauro et al. (2023); Manna and Desai (2024)) have primarily relied on the contribution from unresolved subhalos (without accounting for their central prompt cusps). Therefore, although in this work we neglect the contribution of subhalos beyond the prompt cusps, it is instructive to include a comparison in Fig. 4. Here we show two estimates of the angular distribution of the $J$ factor that result from standard subhalo models. We consider top-level subhalos, subhalos of those subhalos, and so on up to 6 levels of nesting; we verified that further iteration does not change the results. Within each host, subhalos are spatially distributed in accordance with the number density distribution obtained by Ref. Springel et al. (2008). The density profile of each subhalo is NFW, set by the subhalo concentration-mass relation from Ref. Moliné et al. (2017). Finally, the number of subhalos of each mass $m$ is taken to be $\mathrm{d}N/\mathrm{d}m=(A_{\mathrm{sub}}/M)(m/M)^{-\alpha_{\mathrm{sub}}}$ in a host of mass $M$, ranging from a maximum of $m=0.01M$ to a minimum of $m=10^{-6}$ M (approximately the mass of a prompt cusp for dark matter with $m_{\mathrm{DM}}=100$ GeV and $T_{\mathrm{d}}=30$ MeV Delos and White (2023)). We set $(A_{\mathrm{sub}},\alpha_{\mathrm{sub}})=(0.012,2)$ for a more optimistic estimate and $(A_{\mathrm{sub}},\alpha_{\mathrm{sub}})=(0.03,1.9)$ for a less optimistic estimate Sánchez-Conde and Prada (2014). These estimates are similar to the “MAX” and “MED” subhalo models from Ref. Di Mauro et al. (2023) and result in subhalos contributing 8.5 and 83 times the integrated annihilation luminosity of the smooth halo of the Virgo cluster, respectively. In comparison, the cusps contribute integrated luminosity about 200 times that of the smooth halo.
The estimated subhalo contribution shown in Fig. 4 is significantly more spatially extended than the contribution from the cusps. This outcome is due to tidal stripping of the subhalos (e.g. Han et al. (2016)). While prompt cusps are also affected by tidal stripping, their more centrally concentrated structures make them more resistant to tidal effects Penarrubia et al. (2010); Du et al. (2024); Aguirre-Santaella et al. (2025) while also suppressing the impact of any mass loss on the annihilation luminosity Delos and White (2023). At large $\theta$, the more optimistic subhalo estimate lies within a factor of 2 of the cusp contribution, while the less optimistic estimate is lower by at least an order of magnitude at all $\theta$. Note, however, that the subhalo estimates are based on simulations of roughly $10^{5}$ M to $10^{10}$ M subhalos of $10^{12}$ M field halos Diemand et al. (2008); Springel et al. (2008). Calculating the annihilation signal from subhalos involves extrapolating these results by more than 10 orders of magnitude in both subhalo mass and host halo mass (for nested subhalos). Consequently, these estimates are highly uncertain (see also Refs. Okoli et al. (2018); Drakos et al. (2023)). Indeed, a more recent semianalytic subhalo modeling approach found that subhalos contribute less annihilation luminosity than both of the extrapolations we have presented Hiroshima et al. (2018); Ando et al. (2019). It is because of these uncertainties that we do not include the contribution from the subhalos (beyond their central prompt cusps) in our analysis.
II.3 Robustness of the predicted signals
We emphasize that the annihilation signal from the prompt cusps is much clearer than that coming from unresolved subhalos. The initial distribution of cusps emerges directly from the statistics of the cosmological initial conditions Delos et al. (2019a); Delos and White (2023), and tidal stripping during their evolution as substructures can only marginally reduce the annihilation luminosity from the cusps Delos and White (2023). The one extrapolation that we make involves the number of cusps that survive the hierarchical assembly of cosmic structure (as discussed at the end of Section II.1). Reference Delos and White (2023) found that the annihilation rate contributed by surviving cusps rapidly converges, becoming essentially constant by a time that corresponds to $z\simeq 7$ in our scenarios. [^1] Reference Delos and White (2023) found that the annihilation rate is constant by the time that $\sigma_{0}\simeq 3$, where $\sigma_{0}$ is the rms fractional density contrast in linear theory. For dark matter with $m_{\mathrm{DM}}=100$ GeV and $T_{\mathrm{d}}=30$ MeV, that happens at redshift $z\simeq 7$. We assume that this contribution remains constant thereafter. This assumption is well justified, however, because loss of prompt cusps through dynamical friction inspirals should become rare when typical halos are much more massive than the cusps. Cusps can also be disrupted by close encounters with stars, but this effect is only significant near the dense centers of galaxies Stücker et al. (2023a).
There are nevertheless two main modeling uncertainties in the signal from the prompt cusps. First, the numerical factor in Eq. (1), by which cusp coefficients $A$ are evaluated, is uncertain by around 20 percent. We chose the prefactor 24 suggested by Ref. Delos and White (2022), but Refs. Delos et al. (2019a); Ondaro-Mallea et al. (2023) obtained somewhat different prefactors of 26.5 and 19.1, respectively. This 20 percent uncertainty in $A$ translates to about a factor of 1.5 uncertainty in the annihilation rate. Second, Ref. Delos and White (2023) found that the fraction $f_{\mathrm{surv}}$ of the total cusp annihilation luminosity that survives hierarchical assembly is uncertain by about 20 percent. Overall, there is close to a factor of 2 uncertainty in the annihilation signal from the cusps. We follow Ref. Delos et al. (2024) in not incorporating this uncertainty into our results. Since it propagates cleanly through our analyses, the annihilation limits that we derive can be straightforwardly modified to account for any future update to $f_{\mathrm{surv}}$ or the numerical prefactor in Eq. (1). Although this level of uncertainty is significant, we note that conventional annihilation searches using dwarf spheroidal galaxies can face similar or worse levels of uncertainty arising from incomplete knowledge about how the mass is distributed in these systems (e.g. Sanders et al. (2016); Hayashi et al. (2016); Chang and Necib (2021)). The signal from prompt cusps does not face that uncertainty because it depends primarily on the number of cusps, and hence only on the total mass of the dark matter system.
Also, the signal from prompt cusps is sensitive to the cosmological initial conditions on far smaller scales than are constrained by observations. As we discussed in Section II.2, the annihilation rate scales logarithmically with the decoupling temperature $T_{\mathrm{d}}$, since $T_{\mathrm{d}}$ influences the dark matter free-streaming scale. [^2] Reference Delos and White (2023) found that the annihilation rate scales as $\langle\rho_{\mathrm{eff}}\rangle_{0}\simeq 0.08\left[\ln\!\left(\frac{m_{\mathrm{DM}}}{\mathrm{GeV}}\frac{T_{\mathrm{d}}}{\mathrm{GeV}}\right)+36\right]^{5}\bar{\rho}_{0}.$ (14) Our choice to fix $T_{\mathrm{d}}=30$ MeV is mildly conservative (see Section II.2), and dark matter with higher $T_{\mathrm{d}}$ could yield an annihilation signal that is more luminous by a factor of a few. Moreover, the prompt cusp signal is sensitive to the small-scale matter power spectrum more broadly, and the annihilation signal could be greatly altered if either the primordial power spectrum Delos et al. (2018a, b) or the early thermal history of the Universe Delos et al. (2019b, 2023); Ganjoo and Delos (2025); Bae et al. (2025); Fairbairn et al. (2025) deviates from the simplest assumptions. Note that these considerations are not specific to prompt cusps, and they are also relevant to the signal from unresolved subhalos used by many previous annihilation searches (e.g. Ackermann et al. (2015c, b); Liu et al. (2017); Chan and Leung (2017); Bartlett et al. (2022); Di Mauro et al. (2023); Manna and Desai (2024); Cholis and Krommydas (2024)).